Show all steps1111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111111
6. (4 points/1 each) Simplify the following as much as possible without a calculator. Show work! (a) zJust simplify this; you dont have to do the final multiplication.) 99! 97! (b) 8C (c) soCo (d) 26C3 (Simplify similar to part (a).)
Answer #1

a)99!/97! =99*98*97!/97! =99*98=9702

b) 8C7=8!/(7!*(8-7)!)=8!/(7!*1!)=8*7!/7!=8

c)80C0 =80!/(0!*(80-0)!)=80!/(1*80!)=1

d)26C3=26!/(3!*(26-3)!)=26!/(3!*23!)=26*25*24*23!/(3*2*1*23!)=26*25*24/6=26*25*4=2600